OpenStudy (anonymous):

The surface area of a sphere is found using the formula SA=4πr^2, where r is the radius. If the surface area of a sphere is 144π square feet, what is the radius, in feet? A.6 B.12 C.18 D.36 Please Help

OpenStudy (anonymous):

I'm thinking that it's 18 but i'm not sure

OpenStudy (nurali):

4πr^2=144π r^2=144π/4π r^2=36 r=6

OpenStudy (nurali):

However you cannot have a negative radius so we get rid of the negative one and we are left with

OpenStudy (anonymous):

we are left with r=6 right because 144π=4πr2 divide by 4 pi 36=r^2 square root: r=±6 we can't have the negative so we end up with r = 6